OpenStudy (anonymous):

what is the simplified form of... (5/3y-2)-(4/9y^2-4)

OpenStudy (anonymous):

\[\frac{5}{3y-2}-\frac{4}{9y^2-4}\] like that?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

k then we need a common denominator in order to subtract, as with ordinary fractions \[9y^2-4=(3y+2)(3y-2)\] so your common denominator is \[(3y+2)(3y-2)\] subtract as follows \[\frac{5}{3y-2}-\frac{4}{9y^2-4}=\frac{5(3y+2)-4}{(3y+2)(3y-2)}\]then some algebra to clean up the numerator

OpenStudy (anonymous):

you got that part ?

OpenStudy (anonymous):

yah its ether 15y+x/(3y-2)(3y+2) or the bottom are both plus 2

OpenStudy (anonymous):

hold the phone. there is no \(x\) here

OpenStudy (anonymous):

should be \[\frac{15y+6}{(3y+2)(3y-2)}\]